OpenStudy (anonymous):

@ganeshie8 can you help me on the last part of the ques?

OpenStudy (anonymous):

can you write out the steps for me please(:

ganeshie8 (ganeshie8):

familiar with solving the differential equation using separation of variables ?

ganeshie8 (ganeshie8):

\[\dfrac{dy}{dx} = \dfrac{xy}{3}\] this can be rewritten as \[\dfrac{1}{y}dy = \dfrac{x}{3}dx\]

ganeshie8 (ganeshie8):

integrate now

ganeshie8 (ganeshie8):

\[\int \dfrac{1}{y}dy = \int \dfrac{x}{3}dx\]

ganeshie8 (ganeshie8):

can you finish the rest ?

OpenStudy (campbell_st):

so which part is the problem..?

OpenStudy (anonymous):

@perl can you please help? ganeshie left

OpenStudy (perl):

did you get up to the integral part

OpenStudy (anonymous):

its number 3, i just dont get what to do next

OpenStudy (anonymous):

yeah i get it till he left

OpenStudy (perl):

when we integrate we get ln|y| = 1/3* x^2/2 + c

OpenStudy (anonymous):

ohh and then what?

OpenStudy (perl):

then we can raise both sides to the power of e

OpenStudy (perl):

y = e^(x^2/6 + C) do you agree so far?

OpenStudy (anonymous):

yes! so we dont solve for c?

OpenStudy (perl):

we will, but first I want to simplify the equation

OpenStudy (anonymous):

okk

OpenStudy (perl):

ln|y| = 1/3* x^2/2 + C e^ln(y) = e^ ( x^2/6 + C ) y = e^(x^2/6 + C) by law of exponents a^(m+n) = a^m * a^n y = e^(x^2/6) * e^C now bring the e^C in front, call it C again, since e^ constant = constant

OpenStudy (perl):

that gives us the simplified version y = C * e^(x^2/6)

OpenStudy (anonymous):

and thats it?

OpenStudy (anonymous):

thank you!(:

OpenStudy (perl):

now plug in f(0) = 4

OpenStudy (anonymous):

oh hahaha

OpenStudy (perl):

4 = C * e^( 0^2/6) 4 = C * e^0 4 = C * 1

OpenStudy (perl):

and just to be on the safe side, we can confirm with wolfram http://www.wolframalpha.com/input/?i=solve+dy%2Fdx+%3D+xy%2F3+%2C+y%280%29%3D4

OpenStudy (perl):

by the way, can someone tell me why we drop the absolute value bars e^ln|y| becomes y